OpenStudy (anonymous):

what is the remainder when (3x^4 + 2x^3 - x^2 + 2x - 24) / (x + 12)

OpenStudy (anonymous):

I mean x(+2) not (x+12)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (raden):

first, solve for x x+2=0 x = -2 now plugin x=-2 to (3x^4 + 2x^3 - x^2 + 2x - 24) what do you get ?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

I have no idea. What do I do with the exponents

OpenStudy (raden):

(-2)^4 = -2 * -2 * -2 * -2 = 16 (-2)^3 = -2 * -2 * -2 = -8 (-2)^2 = -2 * -2 = 4

OpenStudy (raden):

got it?